How many gram of ammonia should be added to 500mL of 1M ammonium chloride to give a pH 9?(Ka ammonium=5.6x10-10)


1
Expert's answer
2021-12-03T13:52:01-0500

NH3 = 17 g/mole


pH = 9–––> [H+] = 10^-9 M ––––> [OH–] = 10^-5 M


Ka*Kb = Kwater


(5.6×10^-10)*Kb = 10^-14


Kb = 1.8×10^-5 ( Kb of NH3)


[OH–] = nNH3*Kb / nNH4Cl


10^-5 = nNH3*1.8×10^-5 / 0.5


nNH3 = 0.28 mole


0.28*17 = 4.76 g of NH3